OpenStudy (peace420):

Write two like terms using whatever variable(s) you'd like. Be sure to include exponents. Explain why these are like terms.

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (peace420):

nope hah

OpenStudy (skullpatrol):

What is the definition of like or similar terms?

OpenStudy (skullpatrol):

Do you know what a monomial is?

OpenStudy (taylor0402):

Combining like terms means putting all the numbers with the same variables together. In other words, an equation might come to you looking like this: 2x + 3y + 4x = 9 If you looked at the terms, or variables, of the equation, you would see that there are two x terms, 2x and 4x, and they’re being added together (you know this because the 4 has a + in front of it. If it were a – sign, we would know they would be subtracted. Now, you can combine the two like terms, which would look like this: 2x + 4x = 6x Now, you have 6x + 3y = 9. This is as much as you can solve this equation, because you are not given values for either of the variables. Thus, your final answer is 6x + 3y = 9.

OpenStudy (skullpatrol):

That^ does not answer what the question is asking for.

OpenStudy (peace420):

well?

OpenStudy (skullpatrol):

well, what?

OpenStudy (skullpatrol):

What are the answers to the definitions I asked you?

OpenStudy (peace420):

nevermind i figured it out.

OpenStudy (skullpatrol):

Good job :-)

OpenStudy (peace420):

thanks